OpenStudy (anonymous):

Solve x2 − 7x + 12 = 0.

OpenStudy (anonymous):

Can you factor this out?

OpenStudy (anonymous):

x=4,3

OpenStudy (anonymous):

What don't give out the answers!!!

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

oh sorry x2-4x-3x+12 x(x-4)-3(x-4)=0 (x-3)(x-4)=0 either x-4=0 or x-3=0 thus x=4 or 3

OpenStudy (anonymous):

You are still essentially still giving the answer away.

OpenStudy (anonymous):

you can factor this x^2-4x-3x+12=0 x(x-4)-3(x-4)=0 (x-3)(x-4)=0 x-3=0 x-4=0 x=3 x=4

OpenStudy (anonymous):

oh i get it thanks. i have another question

OpenStudy (jiteshmeghwal9):

\[x^2-7x+12=0\]using formula\[-b \pm \sqrt{b^2-4ac} \over 2a\]\[7 \pm \sqrt{49-48}\over2 \]\[7 \pm 1 \over 2\]now as u know there are two distinct roots of a quadratic equation so\[\alpha={7+1\over2}={8 \over 2}=4\]\[\beta={7-1\over2}={6 \over2}=3\]

OpenStudy (jiteshmeghwal9):

so x=4 & x=3.